### v1

Abstract.
This work introduces MultiTRON, an approach that adapts Pareto front approximation techniques to multi-objective session-based recommender systems using a transformer neural network. Our approach optimizes trade-offs between key metrics such as click-through and conversion rates by training on sampled preference vectors. A significant advantage is that after training, a single model can access the entire Pareto front, allowing it to be tailored to meet the specific requirements of different stakeholders by adjusting an additional input vector that weights the objectives. We validate the model’s performance through extensive offline and online evaluation. For broader application and research, the source code [^1] https://github.com/otto-de/MultiTRON is made available. The results confirm the model’s ability to manage multiple recommendation objectives effectively, offering a flexible tool for diverse business needs.
session-based, recommender systems, multi-objective, pareto front
ccs: Applied computing Online shoppingccs: Information systems Recommender systemsccs: Computing methodologies Neural networksccs: Computing methodologies Multi-task learning
1. Introduction
Large e-commerce platforms such as OTTO face the complex task of optimizing diverse revenue streams through personalized recommendations. These systems cater to various business needs, such as sponsored product advertisements, which generate revenue per click, and organic recommendations to maximize conversions. The challenge lies in balancing these goals, as marketing teams prioritize high visibility and click-through rates for sponsored content, whereas sales departments focus on enhancing conversion rates and customer loyalty through organic product listings.
To address these conflicting objectives, multi-objective recommender systems offer a framework capable of optimizing across different metrics within a unified system (Mahapatra and Rajan, 2020; Xie et al., 2021; Jin et al., 2023; Li et al., 2023; Abdollahpouri et al., 2020). This study adapts Pareto front approximation techniques, previously successful in other domains (Ruchte and Grabocka, 2021; Navon et al., 2021; Lin et al., 2019; Tuan et al., 2024), to session-based recommender systems. By training on sampled preference vectors, our model can access the entire Pareto front at inference time, providing an efficient tool to meet the diverse business goals of stakeholders.
2. Methods
Multi-objective session-based recommender systems predict the next item interaction based on prior user activities. Each user session consists of user-item interactions $s_{raw}=[i_{1}^{a_{1}},i_{2}^{a_{2}},\ldots,i_{T}^{a_{T}}]$, where $T$ is the session length, and $i_{t}^{a_{t}}$ represents the action taken on item $i$ at time $t$. Actions include clicking or ordering, typically with orders following clicks. Sessions are modelled as $s:=[(c_{1},o_{1}),(c_{2},o_{2}),\ldots,(c_{T-1},o_{T-1})]$, where $c_{t}$ is the clicked item at time $t$, and $o_{t}$ indicates if the item was ordered in the same session up to time $T$. Our multi-objective recommender model $\mathcal{R}$ leverages past clicks to predict scores $r_{t}^{i}$ for potential item interactions, optimizing the trade-off between click $\mathcal{L}_{c}(c_{t},r_{t}^{i})$ and order $\mathcal{L}_{o}(o_{t},r_{t}^{i})$ losses. Standard scalarization methods  (Lin et al., 2019; Ruchte and Grabocka, 2021) use a fixed preference vector $\pi:=[\pi_{c},\pi_{o}]$, with $\pi_{c}+\pi_{o}=1$, and minimize:
(1) $\mathcal{L}(c_{t},o_{t},\mathcal{R}_{t},\pi)=\pi_{c}\mathcal{L}_{c}(c_{t},\mathcal{R}_{t})+\pi_{o}\mathcal{L}_{o}(o_{t},\mathcal{R}_{t}).$
This approach does not scale for large datasets because each point on the Pareto front requires a separate model. In multi-objective Pareto front approximation, sampling $\pi\sim Dir(\beta)$ from a Dirichlet distribution with parameter $\beta\in\mathbb{R}^{2}_{>0}$ during training and adding it to the input yields a model $\mathcal{M}(\cdot,\pi)$ conditioned on $\pi$ during inference  (Ruchte and Grabocka, 2021; Navon et al., 2021; Dosovitskiy and Djolonga, 2020; Tuan et al., 2024). We adapt this approach to sequential recommender models $\mathcal{R(\cdot,\pi)}$ and conclude from (Dosovitskiy and Djolonga, 2020) that if $\mathcal{R}^{*}$ minimizes:
$\mathbb{E}_{\pi}\mathcal{L}(c_{t},o_{t},\mathcal{R}_{t}(\cdot,\pi),\pi)=\mathbb{E}_{\pi}\left(\sum\nolimits_{k}\pi_{k}\mathcal{L}_{k}(k_{t},\mathcal{R}_{t}(\cdot,\pi))\right)$
then $\mathcal{R}^{*}$ minimizes Equation 1 almost surely w.r.t $\mathbb{P}_{\pi}$. To address the limitation of narrow Pareto fronts (Ruchte and Grabocka, 2021), we also leverage the non-uniformity loss from (Mahapatra and Rajan, 2020) by adding it to the loss:
$\mathcal{L}_{reg}(\pi)=\text{CE}(\hat{\pi}\mid\textbf{1}/2),$
where $\hat{\pi}_{k}:=\pi_{k}\mathcal{L}_{k}/(\pi_{c}\mathcal{L}_{c}+\pi_{o}\mathcal{L}_{o})$, $\textbf{1}/2:=[\frac{1}{2},\frac{1}{2}]$ and CE is the cross-entropy loss. The overall loss is formulated as:
(2) $\mathbb{E}_{\pi}\mathcal{L(\cdot,\pi,\lambda)}=\mathbb{E}_{\pi}\left(\sum\nolimits_{k}\pi_{k}\mathcal{L}_{k}(k_{t},\mathcal{R}_{t}(\cdot,\pi))+\lambda\mathcal{L}_{reg}(\pi)\right),$
with $\lambda\geq 0$ as a regularization parameter. Our model MultiTRON minimizes the loss in Equation 2 to approximate the Pareto front.
3. Experimental Setup
Table 1. Statistics of the datasets used in our experiments.
Diginetica Yoochoose OTTO
Train sessions 187k 7.9M 12.9M
click events 906k 31.6M 194.7M
order events 12k 1.12M 5.1M
Test sessions 18k 15k 1.6M
click events 87k 71k 1.23M
order events 1.1k 1.2k 355k
Items 43k 37k 1.8M
In our experiments we evaluate the proposed approach using three benchmark datasets of varying complexity: Diginetica (DIGINETICA, 2016), Yoochoose (Ben-Shimon et al., 2015), and OTTO (Philipp Normann et al., 2023). These datasets differ in terms of the number of events and the diversity of item sets. Our experiments focus on click and order events, ensuring a minimum item support of five and a session length of at least two clicks for all datasets (Hidasi and Karatzoglou, 2018). We adopt a temporal train/test split approach for training and testing the models. Specifically, the entire last day from the Yoochoose dataset and the entire last week from the Diginetica and OTTO datasets are designated as test datasets (Krichene and Rendle, 2020), with the remainder used for training. Table 1 presents an overview of these datasets. All models are trained on an NVIDIA Tesla V100 GPU with a batch size of 256. MultiTRON uses the session-based Transformer architecture TRON (Wilm et al., 2023), configured with 3 layers and a learning rate of $10^{-4}$. The Dirichlet parameter $\beta=[\frac{1}{2},\frac{1}{2}]$ is fixed, leading to $\pi\sim Dir([\frac{1}{2},\frac{1}{2}])$. The regularization parameter $\lambda$ is tuned between 0.02 and 1.0 for each dataset. For offline evaluation, the Hypervolume Indicator (HV) is employed (Guerreiro et al., 2022), with reference points based on the nadir points (Deb et al., 2010) of each dataset: $r_{D}=[3.86,1.12]$, $r_{Y}=[4.03,0.17]$, $r_{O}=[3.91,1.02]$. The loss functions used are binary cross-entropy loss for the order task ($\mathcal{L}_{o}$) and sampled softmax loss for the click task ($\mathcal{L}_{c}$), as outlined in recent studies (Wilm et al., 2023; Wu et al., 2022).
4. Results
Table 2. Hypervolumes for $\beta=[\frac{1}{2},\frac{1}{2}]$ and different values of $\lambda$ for each dataset. The models are trained on Diginetica for 20 epochs and Yoochoose and OTTO for 10 epochs.
Datasets $\lambda=0.02$ $\lambda=0.2$ $\lambda=0.5$ $\lambda=1$
Diginetica 0.20609 0.20012 0.20605 0.19296
Yoochoose 0.0806 0.0776 0.0838 0.0831
OTTO 1.524 1.537 1.544 1.546
Table 2 presents the results of our offline evaluation. We found that larger values of $\lambda$ led to increased hypervolumes, especially on more complex datasets. The Pareto fronts demonstrating the highest hypervolumes for each dataset are depicted in Figure 1. Importantly, incorporating the sampling parameter $\pi$ into our model does not adversely impact training speed or increase the number of epochs required compared to training models optimized for single objectives with the same architecture. Given the extensive study of the click task in previous research (De Souza Pereira Moreira et al., 2021; Li et al., 2017; Kang and McAuley, 2018; Hidasi and Karatzoglou, 2018; Hidasi et al., 2016; Wilm et al., 2023), we also provide the Recall@20 for $\pi=[1,0]$, which resulted in scores of 0.529 for Diginetica, 0.724 for Yoochoose, and 0.485 for OTTO. Our previous work (Wilm et al., 2023) utilizing the same TRON architecture trained on solely the click task yielded Recall@20 scores of 0.541 ($-2.2\%$), 0.732 ($-1.1\%$), and 0.472 ($+2.8\%$) for these datasets. These results demonstrate that MultiTRON performs comparably to single-objective click task models that utilize the same backbone.
For the online evaluation, we utilized a model trained on OTTO’s private data collected in May 2024. A live A/B test was conducted the following week, with four groups, each assigned different $\pi$ values. The test results confirmed that the offline trade-off between $-\mathcal{L}_{c}$ and $-\mathcal{L}_{o}$ translates into real-world trade-offs between click-through rates (CTR) and conversion rates (CVR). Specifically, higher $-\mathcal{L}_{o}$ values correlated with increased CVR, while higher $-\mathcal{L}_{c}$ values correlated with increased CTR, as detailed in Figure 2.
Figure 1. The best performing Pareto fronts from the offline evaluation on all three datasets showing the trade-off between $\mathcal{L}_{c}$ and $\mathcal{L}_{o}$ for 26 increasing values of $\pi_{o}$.
[…]
[Three line charts]Three line charts showing the Pareto fronts with the best Hypervolume from the offline evaluation. The x-axis shows the click loss $\mathcal{L}_{c}$, and the y-axis shows the order loss $\mathcal{L}_{o}$. The first chart shows the Pareto front of Diginetica ranging from 3.66 to 3.83 in $\mathcal{L}_{c}$ and 0.09 to 0.72 in $\mathcal{L}_{o}$. The second chart shows the Pareto front of Yoochoose ranging from 2.6 to 3.0 in $\mathcal{L}_{c}$ and 0.11 to 0.15 in $\mathcal{L}_{o}$. The third chart shows the Pareto front of OTTO ranging from 2.0 to 3.9 in $\mathcal{L}_{c}$ and 0.18 to 1.02 in $\mathcal{L}_{o}$. Each curve is convex and has similar shape to $f(x)=1/x$ on $\mathbb{R}_{>0}$
Figure 2. Evaluation results from the offline dataset (left) and live A/B test (right). It demonstrates that points on the predicted offline front translate into real-world trade-offs of CTR and CVR. The colored points correspond to differing values of $\pi_{o}$, representing the four different A/B test groups.
[…]
[Two line charts]The evaluation results from the private test dataset (left chart) and online A/B test (right chart). Each plot contains four colored points corresponding to increasing values of $\pi_{o}$, representing the four different A/B test groups. The test results (left) show the negative click loss $-\mathcal{L}_{c}$ on the x-axis and the negative order loss $-\mathcal{L}_{o}$ on the y-axis. The Pareto front on the negative losses range from -2.9 to -2.3 on the x-axis ($-\mathcal{L}_{c}$) and from -0.35 to -0.19 on the y-axis ($-\mathcal{L}_{o}$). The right chart shows the the trade-off of the CTR uplift (in perenct) on the x-axis and the CVR uplift (in perenct) on the y-axis. Values of the CTR uplift (x-axis) range from 0.0 to 9.9 percent and the values of the CVR uplift (y-axis) range from 0.0 to 6.8 percent. Both charts are rapidly decreasing and concave. The positive correlation is observed: higher $-\mathcal{L}_{o}$ correlates with increased CVR, and higher $-\mathcal{L}_{c}$ correlates with enhanced CTR.
5. Conclusion
In this work, we successfully applied a Pareto front approximation technique to multi-objective session-based recommender systems. We demonstrated the practical relevance of MultiTRON through offline evaluation on three benchmark datasets and an online real-world A/B test. The ability of the offline calculated Pareto front to translate into real-world trade-offs of CTR and CVR in a live setting validates the model’s effectiveness in commercial environments.
6. Speaker Bio
Timo Wilm and Philipp Normann are Senior Data Scientists specializing in the design and integration of deep learning models. Felix Stepprath is a Digital Analyst specializing in advanced analytics. All three are members of OTTO’s recommendation team.

### v3

@@
 The results confirm the model’s ability to manage multiple recommendation objectives effectively, offering a flexible tool for diverse business needs.
-session-based, recommender systems, multi-objective, pareto front
-ccs: Applied computing Online shoppingccs: Information systems Recommender systemsccs: Computing methodologies Neural networksccs: Computing methodologies Multi-task learning
+session-based recommender systems, multi-objective, pareto front
+journalyear: 2024copyright: rightsretainedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688048isbn: 979-8-4007-0505-2/24/10ccs: Applied computing Online shoppingccs: Information systems Recommender systemsccs: Computing methodologies Multi-task learningccs: Computing methodologies Neural networks
 1. Introduction
@@
 The challenge lies in balancing these goals, as marketing teams prioritize high visibility and click-through rates for sponsored content, whereas sales departments focus on enhancing conversion rates and customer loyalty through organic product listings.
-To address these conflicting objectives, multi-objective recommender systems offer a framework capable of optimizing across different metrics within a unified system (Mahapatra and Rajan, 2020; Xie et al., 2021; Jin et al., 2023; Li et al., 2023; Abdollahpouri et al., 2020).
-This study adapts Pareto front approximation techniques, previously successful in other domains (Ruchte and Grabocka, 2021; Navon et al., 2021; Lin et al., 2019; Tuan et al., 2024), to session-based recommender systems.
+To address these conflicting objectives, multi-objective recommender systems offer a framework capable of optimizing across different metrics within a unified system (Milojkovic et al., 2020; Mahapatra and Rajan, 2020; Xie et al., 2021; Jin et al., 2023; Li et al., 2023; Abdollahpouri et al., 2020).
+This study adapts Pareto front approximation techniques, previously successful in other domains (Ruchte and Grabocka, 2021; Navon et al., 2021; Lin et al., 2019b; Tuan et al., 2024), to session-based recommender systems.
 By training on sampled preference vectors, our model can access the entire Pareto front at inference time, providing an efficient tool to meet the diverse business goals of stakeholders.
+2. Related Work
+Recent research in multi-objective optimization for recommender systems has focused on balancing different goals such as accuracy, revenue, and fairness (Wu et al., 2023; Ge et al., 2022; Lin et al., 2019a).
+Traditional methods often involve training multiple models, using different loss scalarizations, constraints or initializations (Milojkovic et al., 2020; Rodriguez et al., 2012; Li et al., 2023; Xie et al., 2021).
+However, these approaches become impractical as each point on the Pareto front requires a separate model, leading to high computational costs during training and inference, particularly with large datasets.
+To address these challenges, Pareto Front Learning (PFL) and Pareto Front Approximation (PFA) have emerged as scalable alternatives (Lin et al., 2019b; Hoang et al., 2023).
+These techniques allow a single model to approximate the entire Pareto front, providing flexibility to adjust to different objectives post-training.
+For example, Dosovitskiy and Djolonga (2020) introduced a method that trains a deep neural network on a distribution of losses conditioned by an additional input vector, effectively integrating multiple objectives within a single model.
+Further advancements include Pareto hypernetworks (PHNs), which generate the model parameters based on preference vectors (Navon et al., 2021; Tuan et al., 2024; Hoang et al., 2023).
+Additionally, Exact Pareto Optimal (EPO) Search has been combined with PHNs to ensure convergence to an exact Pareto optimal solution when one exists, or to the closest possible solution otherwise (Mahapatra and Rajan, 2020; Navon et al., 2021).
+This approach requires solving a linear program after each forward pass, leading to slower training.
+Although PHNs are efficient for exploring Pareto fronts, they face scalability challenges when applied to models with extensive parameters, such as those in recommender systems with large item sets.
+To overcome this limitation, Ruchte and Grabocka (2021) proposed incorporating preference vectors directly into the model as input features, thereby eliminating the need for PHNs.
+The Transformer architecture has gained traction in sequential recommender systems, with models like TRON (Wilm et al., 2023) demonstrating strong performance in session-based tasks.
+TRON’s effectiveness in single-objective optimization motivates its extension to multi-objective settings, where its inherent flexibility can be leveraged for Pareto front approximation.
+3. Contributions
+We introduce MultiTRON, a novel extension of Pareto front approximation techniques tailored for multi-objective session-based recommender systems.
+Building on the Transformer-based TRON model (Wilm et al., 2023), MultiTRON leverages sampled preference vectors and a scalarization approach with a customized regularization term to balance competing objectives, such as click-through and conversion rates.
+Our primary contributions include:
+  1. (1)
+    Scalable Pareto Front Approximation: We extend existing Pareto front approximation methods used in other domains by integrating them into a session-based Transformer model, enabling the efficient exploration of trade-offs between multiple objectives without the need for separate models for each point on the Pareto front.
+  2. (2)
+    Regularization for Improved Coverage: We propose a regularization technique, inspired by the non-uniformity loss introduced by Mahapatra and Rajan (2020), which enhances the diversity of solutions along the Pareto front, ensuring better coverage and minimizing the risk of a collapsing front, while maintaining efficient training times.
+  3. (3)
+    Comprehensive Evaluation: MultiTRON is evaluated on a range of RecSys datasets, demonstrating its effectiveness in achieving competitive performance across multiple objectives.
+Additionally, we validate our approach through extensive online A/B testing, confirming the practical applicability in real-world e-commerce environments.
+  4. (4)
+    Open-Source Implementation: To facilitate further research and practical application of Pareto front approximation in session-based recommender systems, we provide an open-source implementation of MultiTRON.
+To the best of our knowledge, MultiTRON is the first model adapting PFA techniques to session-based recommender systems, providing a scalable and flexible solution that meets the diverse needs of modern e-commerce platforms.
 4. Methods
@@
 Actions include clicking or ordering, typically with orders following clicks.
-Sessions are modelled as $s:=[(c_{1},o_{1}),(c_{2},o_{2}),\ldots,(c_{T-1},o_{T-1})]$, where $c_{t}$ is the clicked item at time $t$, and $o_{t}$ indicates if the item was ordered in the same session up to time $T$.
+Sessions are modelled as $s:=[(c_{1},o_{1}),(c_{2},o_{2}),\ldots,(c_{T-1},o_{T-1})]$, where $c_{t}$ is the clicked item at time $t$, and $o_{t}$ indicates if the item was ordered up to time $T$.
+4.1. Recommender Model and Loss Functions
 Our multi-objective recommender model $\mathcal{R}$ leverages past clicks to predict scores $r_{t}^{i}$ for potential item interactions, optimizing the trade-off between click $\mathcal{L}_{c}(c_{t},r_{t}^{i})$ and order $\mathcal{L}_{o}(o_{t},r_{t}^{i})$ losses.
-Standard scalarization methods  (Lin et al., 2019; Ruchte and Grabocka, 2021) use a fixed preference vector $\pi:=[\pi_{c},\pi_{o}]$, with $\pi_{c}+\pi_{o}=1$, and minimize:
+Standard scalarization methods  (Lin et al., 2019b; Ruchte and Grabocka, 2021) use a fixed preference vector $\pi:=[\pi_{c},\pi_{o}]$, with $\pi_{c}+\pi_{o}=1$, and minimize:
 (1) $\mathcal{L}(c_{t},o_{t},\mathcal{R}_{t},\pi)=\pi_{c}\mathcal{L}_{c}(c_{t},\mathcal{R}_{t})+\pi_{o}\mathcal{L}_{o}(o_{t},\mathcal{R}_{t}).$
 This approach does not scale for large datasets because each point on the Pareto front requires a separate model.
-In multi-objective Pareto front approximation, sampling $\pi\sim Dir(\beta)$ from a Dirichlet distribution with parameter $\beta\in\mathbb{R}^{2}_{>0}$ during training and adding it to the input yields a model $\mathcal{M}(\cdot,\pi)$ conditioned on $\pi$ during inference  (Ruchte and Grabocka, 2021; Navon et al., 2021; Dosovitskiy and Djolonga, 2020; Tuan et al., 2024).
-We adapt this approach to sequential recommender models $\mathcal{R(\cdot,\pi)}$ and conclude from (Dosovitskiy and Djolonga, 2020) that if $\mathcal{R}^{*}$ minimizes:
-$\mathbb{E}_{\pi}\mathcal{L}(c_{t},o_{t},\mathcal{R}_{t}(\cdot,\pi),\pi)=\mathbb{E}_{\pi}\left(\sum\nolimits_{k}\pi_{k}\mathcal{L}_{k}(k_{t},\mathcal{R}_{t}(\cdot,\pi))\right)$
+4.2. Pareto Front Approximation
+In Pareto front approximation, sampling $\pi\sim Dir(\beta)$ from a Dirichlet distribution with parameter $\beta\in\mathbb{R}^{2}_{>0}$ during training and adding it to the input yields a model $\mathcal{M}(\cdot,\pi)$ conditioned on $\pi$ during inference  (Ruchte and Grabocka, 2021; Navon et al., 2021; Dosovitskiy and Djolonga, 2020; Tuan et al., 2024).
+We adapt this approach to sequential recommender models $\mathcal{R(\cdot,\pi)}$ and conclude from (Dosovitskiy and Djolonga, 2020) that if $\mathcal{R}^{*}$ minimizes
+(2) $\mathbb{E}_{\pi}\mathcal{L}(c_{t},o_{t},\mathcal{R}_{t}(\cdot,\pi),\pi)=\mathbb{E}_{\pi}\left(\sum_{k\in\{c,o\}}\pi_{k}\mathcal{L}_{k}(k_{t},\mathcal{R}_{t}(\cdot,\pi))\right),$
 then $\mathcal{R}^{*}$ minimizes Equation 1 almost surely w.r.t $\mathbb{P}_{\pi}$.
-To address the limitation of narrow Pareto fronts (Ruchte and Grabocka, 2021), we also leverage the non-uniformity loss from (Mahapatra and Rajan, 2020) by adding it to the loss:
-$\mathcal{L}_{reg}(\pi)=\text{CE}(\hat{\pi}\mid\textbf{1}/2),$
-where $\hat{\pi}_{k}:=\pi_{k}\mathcal{L}_{k}/(\pi_{c}\mathcal{L}_{c}+\pi_{o}\mathcal{L}_{o})$, $\textbf{1}/2:=[\frac{1}{2},\frac{1}{2}]$ and CE is the cross-entropy loss.
+4.3. Regularization and Pareto Front Coverage
+To address the limitation of narrow Pareto fronts (Ruchte and Grabocka, 2021), we also leverage the non-uniformity term from (Mahapatra and Rajan, 2020), defined as:
+(3) $\mathcal{L}_{reg}(\pi)=\text{KL}(g(\hat{\pi})\mid\textbf{1}/2),$
+where $\hat{\pi}_{k}:=\frac{\pi_{k}\mathcal{L}_{k}}{\pi_{c}\mathcal{L}_{c}+\pi_{o}\mathcal{L}_{o}}$, 1/2=[$\textbf{1}/2=\left[\frac{1}{2},\frac{1}{2}\right]$, and KL is the Kullback–Leibler divergence.
+The function $g$ maps $\hat{\pi}$ to a vector of probabilities summing to 1.
+For instance, $g$ could be chosen as the identity or the softmax function.
+By adding this regularization term in Equation 3 to the primary loss function in Equation 2, we avoid the need for solving a linear program after each forward pass, thereby maintaining efficient training speeds.
+This approach yields a Pareto front that approximately intersects with the inverse preference vector $\pi^{-1}=\left[\frac{1}{g(\pi_{c})},\frac{1}{g(\pi_{o})}\right]$ at the point $[\mathcal{L}_{c}^{*}(\cdot,\pi),\mathcal{L}_{o}^{*}(\cdot,\pi)]$ (Mahapatra and Rajan, 2020).
+4.4. Overall Loss Function
 The overall loss is formulated as:
-(2) $\mathbb{E}_{\pi}\mathcal{L(\cdot,\pi,\lambda)}=\mathbb{E}_{\pi}\left(\sum\nolimits_{k}\pi_{k}\mathcal{L}_{k}(k_{t},\mathcal{R}_{t}(\cdot,\pi))+\lambda\mathcal{L}_{reg}(\pi)\right),$
+(4) $\mathbb{E}_{\pi}\mathcal{L(\cdot,\pi,\lambda)}=\mathbb{E}_{\pi}\left(\sum_{k\in\{c,o\}}\pi_{k}\mathcal{L}_{k}(k_{t},\mathcal{R}_{t}(\cdot,\pi))+\lambda\mathcal{L}_{reg}(\pi)\right),$
 with $\lambda\geq 0$ as a regularization parameter. Our model MultiTRON minimizes the loss in Equation 4 to approximate the Pareto front.
@@
 Test sessions 18k 15k 1.6M
-click events 87k 71k 1.23M
+click events 87k 71k 12.3M
 order events 1.1k 1.2k 355k
 Items 43k 37k 1.8M
-In our experiments we evaluate the proposed approach using three benchmark datasets of varying complexity: Diginetica (DIGINETICA, 2016), Yoochoose (Ben-Shimon et al., 2015), and OTTO (Philipp Normann et al., 2023).
+In our experiments, we evaluate the proposed approach using three benchmark datasets of varying complexity: Diginetica (DIGINETICA, 2016), Yoochoose (Ben-Shimon et al., 2015), and OTTO (Philipp Normann et al., 2023).
 These datasets differ in terms of the number of events and the diversity of item sets.
-Our experiments focus on click and order events, ensuring a minimum item support of five and a session length of at least two clicks for all datasets (Hidasi and Karatzoglou, 2018).
+The experiments focus on click and order events, ensuring a minimum item support of five and a session length of at least two clicks for all datasets (Hidasi et al., 2016).
 We adopt a temporal train/test split approach for training and testing the models.
@@
 All models are trained on an NVIDIA Tesla V100 GPU with a batch size of 256.
-MultiTRON uses the session-based Transformer architecture TRON (Wilm et al., 2023), configured with 3 layers and a learning rate of $10^{-4}$.
+MultiTRON uses the session-based Transformer architecture TRON (Wilm et al., 2023), configured with three layers and a learning rate of $10^{-4}$.
+The loss functions used are binary cross-entropy loss for the order task ($\mathcal{L}_{o}$) and sampled softmax loss for the click task ($\mathcal{L}_{c}$(Wilm et al., 2023; Wu et al., 2022).
+We select $g$ as the softmax function because it has smaller gradients than the identity and, in our experiments, provided more stable convergence to the Pareto front.
 The Dirichlet parameter $\beta=[\frac{1}{2},\frac{1}{2}]$ is fixed, leading to $\pi\sim Dir([\frac{1}{2},\frac{1}{2}])$.
@@
 For offline evaluation, the Hypervolume Indicator (HV) is employed (Guerreiro et al., 2022), with reference points based on the nadir points (Deb et al., 2010) of each dataset: $r_{D}=[3.86,1.12]$, $r_{Y}=[4.03,0.17]$, $r_{O}=[3.91,1.02]$.
-The loss functions used are binary cross-entropy loss for the order task ($\mathcal{L}_{o}$) and sampled softmax loss for the click task ($\mathcal{L}_{c}$), as outlined in recent studies (Wilm et al., 2023; Wu et al., 2022).
-4. Results
+6. Evaluation
 Table 2. Hypervolumes for $\beta=[\frac{1}{2},\frac{1}{2}]$ and different values of $\lambda$ for each dataset. The models are trained on Diginetica for 20 epochs and Yoochoose and OTTO for 10 epochs.
@@
 The Pareto fronts demonstrating the highest hypervolumes for each dataset are depicted in Figure 1.
-Importantly, incorporating the sampling parameter $\pi$ into our model does not adversely impact training speed or increase the number of epochs required compared to training models optimized for single objectives with the same architecture.
+Incorporating the sampling parameter $\pi$ into the model does not adversely impact training speed or increase the number of epochs required compared to training models optimized for single objectives with the same architecture.
 Given the extensive study of the click task in previous research (De Souza Pereira Moreira et al., 2021; Li et al., 2017; Kang and McAuley, 2018; Hidasi and Karatzoglou, 2018; Hidasi et al., 2016; Wilm et al., 2023), we also provide the Recall@20 for $\pi=[1,0]$, which resulted in scores of 0.529 for Diginetica, 0.724 for Yoochoose, and 0.485 for OTTO.
 Our previous work (Wilm et al., 2023) utilizing the same TRON architecture trained on solely the click task yielded Recall@20 scores of 0.541 ($-2.2\%$), 0.732 ($-1.1\%$), and 0.472 ($+2.8\%$) for these datasets.
-These results demonstrate that MultiTRON performs comparably to single-objective click task models that utilize the same backbone.
+These results demonstrate that MultiTRON performs comparably to single-objective click task models that use the same backbone.
+Figure 1.
+The best performing Pareto fronts from the offline evaluation on all three datasets showing the trade-off between $\mathcal{L}_{c}$ and $\mathcal{L}_{o}$ for 26 increasing values of $\pi_{o}$.
+[…]
+[Three line charts]Three line charts showing the Pareto fronts with the best Hypervolume from the offline evaluation.
+The x-axis shows the click loss L subscript c, and the y-axis shows the order loss L subscript o.
+The first chart shows the Pareto front of Diginetica ranging from 3.66 to 3.83 in L subscript c and 0.09 to 0.72 in L subscript o.
+The second chart shows the Pareto front of Yoochoose ranging from 2.6 to 3.0 in L subscript c and 0.11 to 0.15 in L subscript o.
+The third chart shows the Pareto front of OTTO ranging from 2.0 to 3.9 in L subscript c and 0.18 to 1.02 in L subscript o.
+Each curve is convex and has similar shape to f(x)= 1/x on R subscript ¿0
 For the online evaluation, we utilized a model trained on OTTO’s private data collected in May 2024.
@@
 Specifically, higher $-\mathcal{L}_{o}$ values correlated with increased CVR, while higher $-\mathcal{L}_{c}$ values correlated with increased CTR, as detailed in Figure 2.
-Figure 1.
-The best performing Pareto fronts from the offline evaluation on all three datasets showing the trade-off between $\mathcal{L}_{c}$ and $\mathcal{L}_{o}$ for 26 increasing values of $\pi_{o}$.
-[…]
-[Three line charts]Three line charts showing the Pareto fronts with the best Hypervolume from the offline evaluation.
-The x-axis shows the click loss $\mathcal{L}_{c}$, and the y-axis shows the order loss $\mathcal{L}_{o}$.
-The first chart shows the Pareto front of Diginetica ranging from 3.66 to 3.83 in $\mathcal{L}_{c}$ and 0.09 to 0.72 in $\mathcal{L}_{o}$.
-The second chart shows the Pareto front of Yoochoose ranging from 2.6 to 3.0 in $\mathcal{L}_{c}$ and 0.11 to 0.15 in $\mathcal{L}_{o}$.
-The third chart shows the Pareto front of OTTO ranging from 2.0 to 3.9 in $\mathcal{L}_{c}$ and 0.18 to 1.02 in $\mathcal{L}_{o}$.
-Each curve is convex and has similar shape to $f(x)=1/x$ on $\mathbb{R}_{>0}$
 Figure 2.
@@
 [Two line charts]The evaluation results from the private test dataset (left chart) and online A/B test (right chart).
-Each plot contains four colored points corresponding to increasing values of $\pi_{o}$, representing the four different A/B test groups.
-The test results (left) show the negative click loss $-\mathcal{L}_{c}$ on the x-axis and the negative order loss $-\mathcal{L}_{o}$ on the y-axis.
-The Pareto front on the negative losses range from -2.9 to -2.3 on the x-axis ($-\mathcal{L}_{c}$) and from -0.35 to -0.19 on the y-axis ($-\mathcal{L}_{o}$).
-The right chart shows the the trade-off of the CTR uplift (in perenct) on the x-axis and the CVR uplift (in perenct) on the y-axis.
-Values of the CTR uplift (x-axis) range from 0.0 to 9.9 percent and the values of the CVR uplift (y-axis) range from 0.0 to 6.8 percent.
+Each plot contains four colored points corresponding to increasing values of pi subscript o, representing the four different A/B test groups.
+The test results (left) show the negative click loss - L subscript c on the x-axis and the negative order loss - L subscript o on the y-axis.
+The Pareto front on the negative losses range from -2.9 to -2.3 on the x-axis (- L subscript c) and from -0.35 to -0.19 on the y-axis (- L subscript o).
+The right chart shows the the trade-off of the CTR uplift (in percent) on the x-axis and the CVR uplift (in percent) on the y-axis.
+Values of the CTR uplift (x-axis) range from 0.0 to 9.9, and the values of the CVR uplift (y-axis) range from 0.0 to 6.8 percent.
 Both charts are rapidly decreasing and concave.
-The positive correlation is observed: higher $-\mathcal{L}_{o}$ correlates with increased CVR, and higher $-\mathcal{L}_{c}$ correlates with enhanced CTR.
+The positive correlation is observed: higher - L subscript o correlates with increased CVR, and higher - L subscript c correlates with enhanced CTR.
 7. Conclusion
 In this work, we successfully applied a Pareto front approximation technique to multi-objective session-based recommender systems.
-We demonstrated the practical relevance of MultiTRON through offline evaluation on three benchmark datasets and an online real-world A/B test.
+MultiTRON enables a single model to access the entire Pareto front, offering a scalable solution for balancing competing objectives without the need for multiple models.
+We also introduce a regularization term, to improve Pareto front coverage and convergence.
+The practical relevance of MultiTRON was demonstrated through offline evaluation on three benchmark datasets, as well as an online real-world A/B test.
 The ability of the offline calculated Pareto front to translate into real-world trade-offs of CTR and CVR in a live setting validates the model’s effectiveness in commercial environments.